ABC, Inc is planning the purchase of a new equipment which will cost $34,485. The project is expected to last for 6 years. The equipment will have a book value of $3,336 at the end of Year 6. The increase in net working capital is expected to be $2,231, all of which will be recouped at the end of the project. The project is expected to have annual operating cash flows of $12,405. What is the Total Cash Flow in Year 6 of the project if the equipment can be sold for $5,163 and the tax rate is 25%?

Note: In the last year of the project, the Total Cash Flow = Operating Cash Flow + Terminal Cash Flow

Please show the steps you used to answer this as I am trying to learn this.
Answer #1

Answer:

Year 6 Cash flows:

Operating Cash Flow:

Annual operating cash flows = $12,405

Terminal Cash Flow:

Book value of the equipment = $3,336

Salvage value of the equipment = $5,163

Gain = 5163 - 3336 =1,827

Tax on Gain =1827 * 25% = $456.75

Salvage value net of tax = 5163 - 456.75 = 4,706.25

Recovery of net working capital = $2,231

Terminal cash flow = Salvage value net of tax + Recovery of net working capital = 4706.25 + 2231 = 6,937.25

Year 6 Total Cash Flow = Operating Cash Flow + Terminal Cash Flow = $12,405 + 6,937.25 = $19,342.25

Total Cash Flow in Year 6 of the project = $19,342.25